OpenStudy (linn99123):

Adam goes to a skating rink on Tuesdays for $8.50. While there, he pays $1.50 for each item he buys at the concession stand. Which equation best represents this information? y = 8.50x + 1.50 y = 8.50x - 1.50 y = 1.50x - 8.50 y = 1.50x + 8.50

OpenStudy (bibby):

1.50 for each. meaning: x=1 -> 8.50 (rink cost) + 1.50 x=2 -> 8.50 (rink cost) + 3 what do you think

OpenStudy (linn99123):

I dont understand what that is bibby /.\

OpenStudy (linn99123):

like all that stuff u wrote is just numbers and letters and signs

OpenStudy (bibby):

you get charged 1.50 for each thing you buy that's multiplication

OpenStudy (linn99123):

i think i should just fail this all so they will throw me out then when im in school thell throw me out there its a win win sitation for me

OpenStudy (bibby):

D y = 1.50x + 8.50 1.50 for each thing he buys plus the constant cost of 8.50 to get in

OpenStudy (anonymous):

The x represents how much items he bought.

OpenStudy (anonymous):

So lets say, Adam bought 5 items for $1.50. y = 1.50(5) + 8.50 y= 7.50 + 8.50 y = $16.00